You have just invested 100,000 from your Great Aunt. Rather than spend the money you are going to invest the money and use the annual interest to pay for private schooling. The annual tuition is $5150. You have decided to invest the money in two different accounts one paying 7% the other paying 4%/ How much should you invest in each account to earn as close to $5150 as possible
Answer by Boreal(15235)

7% account=x
4% account=100,000-x
interest earned is .07x + .04(100,000-x)=.07x+4000-0.04x=0.03+4000
That is to equal $5150
therefore, 0.03x=$1150
x=$38,333 which at 7% yields $2683.33
100000-x=61,666.67 which at 4% yields $2466.67
Those add to $5150
$2683.33 at 7%
$2466.67 at 4%